### v2

# Disentangling the Impact of Quasiparticles and Two-Level Systems on the Statistics of Superconducting Qubit Lifetime

Abstract
Temporal fluctuations in the superconducting qubit lifetime, $T_{1}$, bring about additional challenges in building a fault-tolerant quantum computer. Although exact mechanisms remain unclear, $T_{1}$ fluctuations are generally attributed to the strong coupling between the qubit and a few near-resonant two-level systems (TLSs) that can exchange energy with an assemble of thermally fluctuating two-level fluctuators (TLFs) at low frequencies. Here, we report $T_{1}$ measurements of qubits with different geometrical footprints and surface dielectrics as a function of temperature. By analyzing the noise spectrum of the qubit depolarization rate, $\Gamma_{1}=1/T_{1}$, we can disentangle the impact of TLSs, non-equilibrium quasiparticles (QPs), and equilibrium (thermally excited) QPs on the variance in $\Gamma_{1}$. We find that $\Gamma_{1}$ variances in the qubit with a small footprint are more susceptible to the QP and TLS fluctuations than those in the large-footprint qubits. Furthermore, the QP-induced variances in all qubits are consistent with the theoretical framework of QP diffusion and fluctuation. We suggest that these findings can offer valuable insight for future qubit design and engineering optimization.
Superconducting qubits are a promising candidate for building a fault-tolerant quantum computer due to their flexibility in circuit design [1, 2, 3, 4, 5]. This progress has been significantly propelled by the enhancement of the qubit lifetime, $T_{1}$, from the nanosecond to millisecond range over the past decades, advanced by improvements in material science, nanofabrication technology, and microwave design and engineering [6, 7, 8]. Meanwhile, it has been observed that the lifetime $T_{1}$, as well as the decoherence time $T_{2}$ and the qubit frequency $\omega_{\text{q}}$, can fluctuate across the time and frequency domains; the qubits with higher $T_{1}$ show larger fluctuations [9, 10, 11, 12, 13]. Since the gate fidelity is primarily limited by qubit coherence, these fluctuations will introduce additional challenges and obstacles to quantum error correction and future scalability.
Although the underlying physics is not yet fully understood, two-level defects residing in amorphous dielectrics at material surfaces are suggested as the primary source of decoherence in superconducting quantum devices [14, 15, 16, 17, 18]. High-frequency, near-resonance TLSs strongly couple to the qubit and dissipate energy to an environment at ambient temperature, causing qubit depolarization. Recently, Müller et al. pointed out that high-frequency TLSs can also interact with thermally fluctuating TLFs with $\hbar\omega\ll k_{\text{B}}T$ through dipole-dipole interactions, therefore fluctuating the qubit $T_{1}$ [9]. Subsequent experimental studies of both flux-tunable and fixed-frequency transmon qubits support this theory by showing the signatures of a few TLSs responsible for the qubit lifetime fluctuation. Specifically, the spectral density of the qubit lifetime extracted from the measurements exhibit Lorentzian and $1/f$-type behavior [10, 12, 11].
Since the number of QPs can fluctuate due to mechanisms such as generation and recombination [19], it is natural to hypothesize that these exact mechanisms can lead to fluctuations in the coherence of superconducting qubits. To elucidate the observed power law relationship between the average and standard deviation of the qubit $T_{1}$, the assumption that the qubit $T_{1}$ is limited by the population of thermally excited QPs was made for this purpose [20, 21]. The same assumption was used to explain the temporal fluctuations of the qubit $T_{1}$ extracted from the double exponential decay in the relaxation curves [22]. By examining the $T_{1}$ fluctuations at a few selected temperatures, Müller et al. briefly discussed the contributions of both TLS and QPs to the $T_{1}$ fluctuations in the same qubit and suggested that if there is any, the fluctuation due to QPs is likely to be smaller than the one associated with strongly coupled TLSs [9].
Although many studies have focused on understanding the qubit $T_{1}$ fluctuations induced by TLS or QP, it is interesting to investigate and disentangle both fluctuating sources in the same device to better understand the decoherence mechanisms and propose strategies for mitigating these noise channels. In this work, we perform $T_{1}$ measurements on three fixed-frequency transmon qubits with different geometrical footprints and surface dielectrics over a temperature range of 7-153 mK. Continued $T_{1}$ measurements of $\sim$72 hours at each temperature allow us to analyze the spectral density of qubit fluctuations in terms of the depolarization rate, $\Gamma_{1}$, across a broad range of frequencies, from which we can extract the contributions of the TLS and QPs to the variance in $\Gamma_{1}$. We further analyze the QP densities near the qubit junction based on a diffusion model and provide quantitative support on the experimental observations of qubit $\Gamma_{1}$ variances induced by non-equilibrium and equilibrium QPs, respectively.
We label the three qubits under study as A, B, and C, all fabricated on two separate but identical Sapphire chips (A and B are on one chip and C on another). Both chips share the same qubit layout, and each qubit has its own $\lambda/4$ coplanar waveguide resonator that is inductively coupled to a common feedline for qubit dispersive readout. Qubits A and B use a 160 nm Nb thin film for the patterned capacitor pads, while qubit C has additionally in situ sputtered $\sim$10 nm Ta encapsulated on the top of the Nb film. Qubit A has small pads (120$\times$510 $\mu\text{m}^{2}$) separated by a 20 $\mu$m gap, and qubits B and C adopt the same geometry, specifically, 150$\times$720 $\mu\text{m}^{2}$ pads with a 150 $\mu$m gap [23]. The fabrication processes are kept nominally identical for the two chips, and the exact size of Al/AlOx/Al junction is used for all qubits. The two chips are separately packaged in the same type of enclosures (Cu coated with Au), wired on two separate measurement chains with nominally identical configurations, and characterized in the same cooldown. Additional details on qubit fabrication, material characterization, and microwave measurements can be found in Ref. 13. Regarding TLS losses, our microwave simulation based on a 3D model shows that the energy participation ratio (EPR) of the surface dielectric in qubit A is about two times larger than in qubits B and C [23]. The dielectric losses of the surface oxide on metal pads of qubits A and B are the same, reported to be on the order of $\delta_{\text{NbO}_{x}}\sim 10^{-2}$ [24, 25]; the dielectric loss $\delta_{\text{TaO}_{x}}\sim 10^{-3}$ [26] of qubit C is about an order of magnitude lower than that of qubits A and B.
Figure 1: (a) Typical $\Gamma_{1}$ temporal fluctuations of qubit B at 7 mK (blue) and 153 mK (red), respectively. At low temperatures, typical telegraphic noise of $\Gamma_{1}$ fluctuations with instantaneous switching and finite dwell time exhibits signatures of the TLS-TLF interaction. At high temperatures, the white noise behavior of $\Gamma_{1}$ indicates that the QP fluctuations are the dominating noise source. (b) The noise spectrum of qubit depolarization rate $S_{\Gamma_{1}}$ at 7 and 153 mK are plotted as the solid symbols. The two data sets are fitted with equation $S_{\Gamma_{1}}(f)=a/f+b$ and shown as the solid lines. The dashed lines are the fitted curves at all other temperatures, illustrating the transition of the dominating noise source from TLSs at low temperatures to QPs at high temperatures.
At each elevated temperature, we wait two hours to thermalize the qubits before performing qubit $T_{1}$ measurements consecutively for about 72 hours. The measurements of qubits A and B are interleaved, while qubit C’s are continuous. Each $\Gamma_{1}$ point in Fig. 1 is extracted by fitting a single-exponential decay of the qubit energy-relaxation curve that typically takes 6 to 10 minutes, depending on the lifetime length of each qubit. In the following analyses, we assume that the TLSs and QPs primarily determine the depolarization rate of our qubits due to the following facts: (1) the Purcell rate of the qubit is experimentally verified to be below 1 kHz, (2) the loss due to trapped flux in the pads is negligible because all qubits are fixed-frequency transmons with a single, sub-micron Josephson junction, and (3) the enclosure is designed to have the lowest resonance mode at 10 GHz, well above the qubit frequency (approximately 4 GHz). Therefore, we can write the qubit depolarization rate as the sum of the TLS and QP losses, i.e., $\Gamma_{1}=\Gamma_{\text{TLS}}+\Gamma_{\text{QP}}$, with $\Gamma_{\text{TLS}}=p\delta_{\text{TLS}}$. Here $p$ is the geometry-dependent EPR, and $\delta_{\text{TLS}}$ is determined by the surface dielectric of the metal pads. In this work, we only consider the dielectric loss from the metal-air interface [13].
Figure 1(a) shows the temporal $\Gamma_{1}$ fluctuations of qubit B at 7 mK (blue) and 153 mK (red), respectively. At low temperatures, the $\Gamma_{1}$ fluctuation resembles similar features as reported in  [9, 10, 11, 12], exhibiting signatures of the TLS-TLF interaction. At higher temperatures, the $\Gamma_{1}$ fluctuation shows a white noise behavior, consistent with the behavior resulting from fluctuations in the QP number [19]. To disentangle the contributions of TLSs and QPs to the $\Gamma_{1}$ fluctuations, we convert the temporal fluctuations to spectral density by performing the Fourier transform of the autocorrelation function of the qubit depolarization rate $\langle\Gamma_{1}(t)\Gamma_{1}(0)\rangle$, following the method reported in Ref. 9. The obtained spectral densities of qubit B at 7 mK (blue dot) and 153 mK (red dot) are shown in Fig. 1(b) as a function of frequency. At low temperatures (e.g., $T=7$ mK), we observe a clear $1/f$ behavior at low frequency and a white noise behavior at high frequency. The former behavior is predicted within the interacting TLS-TLF model, where a few TLSs undergo frequency fluctuations due to interactions with an assembly of thermal TLFs [9]. On the other hand, it is known that the number of QPs can fluctuate due to the dynamics of QP generation and recombination. This results in a Lorentzian-type spectral density [27, 19]. Since the frequencies we measured here are much smaller than the characteristic QP recombination rate ($\sim$ kHz), the Lorentzian profile can be well approximated as constant white noise. Therefore, we fit the data with the equation $S_{\Gamma_{1}}(f)=a/f+b$, with results shown as thick solid lines in Fig. 1(b). At higher temperatures (e.g., $T=153$ mK), the spectral density exhibits mainly white noise behavior, indicating that QPs dominate the fluctuation. In Fig. 1(b), we also plot the fitted curves at different temperatures to illustrate the transition of the dominant noise source from TLSs at low temperatures to QPs at high temperatures.
Figure 2: Temperature dependence of the average depolarization rate $\mu(\Gamma_{1})$ for all three qubits. Solid symbols represent the data, dash-dotted lines are the fitted contribution from TLS, and dashed lines are the fitted contribution from non-equilibrium and equilibrium QPs. Solid lines show the sum of the TLS and QP losses, which is in agreement with the measured data. The extracted fitting parameters are provided in Table 1.
We next analyze the impact of QP and TLS losses on the qubit average $\Gamma_{1}$, $\mu(\Gamma_{1})$, as a function of temperature. The measurement data is shown as solid symbols in Fig. 2. It is seen that $\mu(\Gamma_{1})$ of all three qubits show weak temperature dependence when temperatures are below $\sim$100 mK, then exponentially increase and approach the same value around 153 mK. Based on established TLS and QP theories, we can unravel the contributions of TLS and QP loss to $\mu(\Gamma_{1})$. First, although both the relaxation and excitation rate of the qubit due to a single TLS depend on temperature, the sum of both, i.e., the total TLS depolarization rate contributing to the qubit is temperature independent [28, 29]. Secondly, the contributions of QPs to the qubit $\Gamma_{1}$ can be written as [19, 30, 31]:
$\Gamma_{\text{QP}}=\frac{16E_{\text{J}}}{\hbar\pi}\sqrt{\frac{E_{\text{C}}}{8E_{\text{J}}}}\sqrt{\frac{2\Delta}{\pi k_{\text{B}}T}}\exp\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)K_{0}\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)x_{\text{QP}},$ (1)
[…]
From the fits, all three qubits show similar gap energy $\Delta/2\pi\sim 40$ GHz, which is consistent with the values reported in the literature [33, 34], although slightly lower. This is likely due to differences in Al film thicknesses and fabrication processes. The extracted $\Gamma_{\text{TLS}}$ values are shown as dash-dotted lines in Fig. 2, and the improvement of $\Gamma_{\text{TLS}}$ from qubit A to C can be explained by either the simulated EPR or the differences in dielectric loss. The ratio of $\Gamma_{\text{TLS}}$ between qubits A and B is approximately 1.9, which matches our EPR simulation; the ratio between qubit B and C is about 3.9, smaller than the reported $\delta_{\text{NbO}_{x}}/\delta_{\text{TaO}_{x}}\sim 10$ [26, 25, 24]. This discrepancy might arise because the Ta thin film in qubit C encapsulates only the top surface of Nb but not the sidewall [13]. The contribution from QPs is shown as dashed lines in Fig. 2. As expected, $\Gamma_{\text{QP}}$ is dominated by $x^{\text{th}}_{\text{QP}}$ at temperatures, $k_{\text{B}}T$, above the Al gap, $\Delta$, and by $x^{0}_{\text{QP}}$ when $k_{\text{B}}T$ is below $\Delta$. However, it should be noted that although the junctions in all three qubits are nominally identical, the fitted $x^{0}_{\text{QP}}$ shows geometric dependence, with $x^{0}_{\text{QP}}$ ratio $\sim 2.5$ between small pads (qubit A) and large pads (qubits B and C). We prove that this difference can be qualitatively explained in the following, with more details in the Supplementary Material [23].
We adopt a phenomenological energy-dependent diffusion model [35, 36, 37] to calculate the QP density arriving at the junction, assuming that the non-equilibrium QPs generated by high-energy radiation distribute uniformly in the qubit pads and elastically diffuse to the junction. The incoming QP energy is set to be near the gap energy of the system, and Neumann boundary conditions in the substrate-pad interface and junction are flux continuity. In the non-equilibrium regime, $x^{0}_{\text{QP}}$ in qubit A is about $2.7$ times larger than the ones in qubits B and C due to the shorter distance that QPs travel in small pads to the junction. This value is close to the ratio obtained from experimental fitting. In the equilibrium regime, QPs are dominated by those thermally generated from the low-gap Al (instead of high-gap Nb). Since all qubits studied in this experiment share the same design of Al/AlOx/Al Josephson junction, we expect $x^{\text{th}}_{\text{QP}}$ to be the same.
Figure 3: Temperature dependence of qubit $\Gamma_{1}$ variance contributed by (a) QPs and (b) TLSs. In Figure (a), solid symbols represent the data extracted from experiments, and solid lines are the fitting based on Eq. (2). The extracted fitting parameters are provided in Table 1.
To quantify the contributions of QPs and TLSs to the variance of the qubit relaxation rate $\Gamma_{1}$, we integrate the white noise and $1/f$ noise spectral densities over the measured frequency range at each temperature [see the fitted curves shown in Fig. 1(b)], which allows us to separately determine the variances of the qubit depolarization rates induced by QPs, $\sigma^{2}_{\text{QP}}(\Gamma_{1})$, and TLSs, $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$. The results are plotted as solid symbols in Fig. 3(a) and (b).
We first discuss the impact of QP fluctuations on the $\Gamma_{1}$ variance. From Fig. 3(a), we can see that $\sigma^{2}_{\text{QP}}(\Gamma_{1})$ in all qubits exhibit a very similar trend: a weak temperature-dependence at $T<100$ mK followed by a rapid increase at higher temperatures. It is also seen that in the low-temperature regime, qubit A is much more susceptible to the QP fluctuations, with $\sigma^{2}_{\text{QP}}(\Gamma_{1})$ at least an order of magnitude higher than those observed for qubits B and C, while the latter two have similar $\Gamma_{1}$ variances. With the QP diffusion and fluctuation model, we can provide quantitative support in understanding these observations and further disentangle the impacts of non-equilibrium and thermal QPs on the qubit fluctuations. Assuming the number of QPs follows a Poisson distribution [19], we have the QP variance $\sigma^{2}(N_{\text{QP}})$ equals its mean value $\mu(N_{\text{QP}})$. Therefore, the variance $\sigma^{2}_{\text{QP}}(\Gamma_{1})$ is given by [23]:
$\sigma^{2}_{\text{QP}}(\Gamma_{1})=\eta^{2}(T)\left(\frac{x_{\text{QP}}^{0}}{n_{\text{CP}}V^{0}_{\text{eff}}}+\frac{x_{\text{QP}}^{\text{th}}}{n_{\text{CP}}V^{\text{th}}_{\text{eff}}}\right),$ (2)
with $\eta(T)=\frac{16E_{\text{J}}}{\hbar\pi}\sqrt{\frac{E_{\text{C}}}{8E_{\text{J}}}}\sqrt{\frac{2\Delta}{\pi k_{\text{B}}T}}\exp\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)K_{0}\left(\frac{\hbar\omega_{\text{q}}}{2k_{\text{B}}T}\right)$. Here, the non-equilibrium and equilibrium QPs are assumed to be independent [30]. In Eq. (2), we define $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ as the effective volumes for non-equilibrium and thermal QPs, respectively, to reflect their distinct contributions to the fluctuations. Using previously obtained $x^{0}_{\text{QP}}$, $x^{\text{th}}_{\text{QP}}$, and $\Delta$ [see Table 1], we can fit the experimentally extracted $\sigma^{2}_{\text{QP}}(\Gamma_{1})$ and obtain the values of $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ for each qubit. The fitted curves are shown as solid lines in Fig. 3(a), and the fitted $V^{\text{0}}_{\text{eff}}$ and $V^{\text{th}}_{\text{eff}}$ are also listed in Table 1. In the non-equilibrium regime, the effective volume is expected to be proportional to $(v_{\text{F}}\cdot\tau_{\text{r}})^{3}$, with the Fermi velocity $v_{\text{F}}$ being the same for all qubits, and $\tau_{\text{r}}$ the QP recombination rate that is inversely proportional to $x_{\text{QP}}^{0}$ [19]. Therefore, we have $V^{0}_{\text{eff}}\propto(1/x_{\text{QP}}^{0})^{3}$. We notice that the ratio of $V^{0}_{\text{eff}}$ between device C (B) and A is about 13 (5) from the fitting. Indeed, this aligns with our data: the ratio of $x_{\text{QP}}^{0}$ between device A and C (B) is about 2.5 ($2.5^{3}\sim 15$). Moreover, the difference in $x_{\text{QP}}^{0}$ can be supported by the diffusion process, from which the $x_{\text{QP}}^{0}$ ratio is calculated to be 2.7 between different pad sizes. In the thermal regime, the QP density dominated by the thermal generation in the Al junction leads increases significantly but does not substantially suppress the $V_{\text{eff}}^{\text{th}}$. On the other hand, $V_{\text{eff}}^{\text{th}}$ cannot exceed the upper limit set by the non-equilibrium QPs from the qubit pads when $k_{\text{B}}T$ approaches the Al gap energy. Therefore we approximate the $V_{\text{eff}}^{\text{th}}$ as a constant in Eq. (2). From the fit, we find the effective volume $V^{\text{th}}_{\text{eff}}$ for all the qubits approaches the same value, close to the volume of the Josephson junction $\sim 0.013\,\mu\textrm{m}^{3}$, which is as expected since all qubits consist of the same Al junctions. Here, we define the junction volume as the product of the junction area and the total thickness of Al layers.
Table 1: Fitting parameters extracted for three different transmon qubits. Here, $x_{\text{QP}}^{0}$ is the density of non-equilibrium QPs, $\Delta$ is the superconducting gap energy of Al, $V_{\text{eff}}^{0}$ and $V_{\text{eff}}^{\text{th}}$ are the effective volumes for non-equilibrium and thermal QPs, and $\Gamma_{\text{TLS}}$ is the depolarization rate from TLS.
Qubit $x_{\text{QP}}^{0}$ $\Delta/2\pi$ (GHz) $V_{\text{eff}}^{0}$ ($\mu\text{m}^{3}$) $V_{\text{eff}}^{\text{th}}$ ($\mu\text{m}^{3}$) $\Gamma_{\text{TLS}}$ ($\mu\text{s}^{-1}$)
A $1.4\times 10^{-7}$ 38.0 0.062 0.025 1.2×$\times 10^{-2}$
B $5.5\times 10^{-8}$ 38.2 0.290 0.039 6.2×$\times 10^{-3}$
C $5.5\times 10^{-8}$ 39.6 0.807 0.037 1.6×$\times 10^{-3}$
In Fig. 3(b), we plot the $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ induced by TLSs. Compared to the variance from QPs, the $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ here show an overall weak dependence on temperature, with some local features. It is also observed that $\sigma^{2}_{\text{TLS}}(\Gamma_{1})$ of qubit A is notably larger than qubits B and C, while the latter two show some consistency in the temperature range. We suggest that this can be qualitatively explained by TLS theory as well. The variance of qubit depolarization due to the interaction between a single TLS and an assemble of TLFs is given by [29]:
$\sigma^{2}_{\text{TLS}}(\Gamma_{1})=A^{4}\frac{16\gamma_{2}^{2}\omega_{\delta}^{2}}{(\gamma_{2}^{2}+\omega_{\delta}^{2})^{4}}\left[\sum_{i}g_{i}^{2}\left(1-\tanh^{2}\frac{\omega_{\text{t,i}}}{2k_{\text{B}}T}\right)\right]^{2},$ (3)
where $A$ denotes a constant related to the coupling strength and qubit matrix element, $\gamma_{2}$ describes the linewidth of the TLS, and $g_{i}$ the coupling strength between the TLS and the $i$th TLF, with frequency $\omega_{\text{t,i}}$. When the temperature increases, the TLS linewidth gets larger, suppressing the amplitude of the Lorentzian and, thus, the variance of the fluctuation. On the other hand, higher temperature leads to more frequent switching of the TLFs, increasing the last factor in Eq. (3). This discussion is based on a single TLS. In practice, multiple TLSs can couple simultaneously to the qubit, each with different dephasing rates and frequencies. As a result, the above two competing factors for different TLSs may lead to the observed behavior in Fig. 3(b) [23].
Our work underscores a practical need to stabilize qubit lifetimes. On one hand, theoretical proposals based on driving the qubits [38] or the TLSs [29] could provide ways to stabilize the fluctuations due to TLSs without affecting the average lifetime. However, protocols to reduce QP fluctuations have yet to be developed, but may benefit from strategies to reduce QPs in general [36, 39, 40, 34].
We perform prolonged, statistical $T_{1}$ measurements on three superconducting transmon qubits across a broad temperature range. Different geometrical footprints and superconducting materials of these qubits enable us to disentangle the impacts of QPs and TLSs on the average and fluctuation of the qubit lifetime. The improvements in the average qubit lifetime can be explained by an optimization of the qubit design as well as a reduced loss of surface dielectric. More importantly, the QP densities near the Josephson junction depend on the qubit footprints, i.e., higher density in the small-footprint qubit. We further propose the concept of effective QP volume, which scales cubically with the inverse of the non-equilibrium QP density and provides quantitative support for understanding the discrepancy between the lifetime fluctuations in qubits with different footprints. We suggest that these findings can guide future qubit design and engineering optimization.
Acknowledgements.
This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359.